Assume a wheat producing farmer engaging in selling its product under perfect competition market faces cost functions as TC= Q3 -2Q2 +8Q and Average revenue of the farmer is given as Birr 8 . Having this information, A) Determine the optimal level of output and price in the short run. B) Calculate the economic profit (loss) the farmer will obtain (incur


1
Expert's answer
2021-12-29T12:30:39-0500

Solution:

A.). Optimal level of output is where MR = MC:

Under perfect competition = AR = MR = 8

MC = "\frac{\partial TC} {\partial Q}" = 3Q2 – 4Q + 8

8 = 3Q2 – 4Q + 8

Q = 1.33

The optimal level of output in the short-run is = 1.33

Price = 8

 

B.). Economic profit = TR – TC

TR = P "\times" Q = 8 "\times" 1.33 = 10.64

TC = Q3 -2Q2 +8Q = 1.333 – 2(1.332) + 8(1.33) = 2.353 – 3.538 + 10.64 = 9.46

Economic Profit = 10.64 – 9.46 = 1.18